Grade 12th passGeneral Physics

When quantity of electricity equal to that required to liberate 2.24 L hydrogen from 0.1M aqueous H2SO4 is passed then mass of copper that will be deposited at cathode in electrolysis of0.2M solution of Copper sulphate?

Parul Rajput
9 Years agoGrade 12th pass

1 Answer

Vikas TU
ApprovedApproved Tutor Answer9 Years ago
From the Farday’s law of electrolusis we know that mass deposited is given as:
m = (Q/F)*(M/z)
or
m/M =moles =Qz/F
From M1V1=M2V2
2.24 *0.1 => 0.224 moles of CuSO4 required.
Hence
0.224 =Q*2/F
Q = 0.112F => 10800 Coulamb of charge would be required.